# Strong Gravitational Lensing Effects of the Rotating Short-Haired Black Hole and Constraints from EHT Observations

### v2

@@ III Strong Gravitational Lensing in the Spacetime of the Rotating Short-hairy Black Hole
$a$ $Q_{m}$ $\bar{a}$ $\bar{b}$ $u_{m}/R_{s}$ $u_{m}/R_{s}$
$k=1$ $k=1.5$ $k=1$ $k=1.5$ $k=1$ $k=1.5$
-0.2 0 0.8796 0.8796 -0.2810 -0.2810 2.9788 2.9788
0.3 0.9048 0.8911 -0.2570 -0.2824 2.8331 2.9543
0.4 0.9354 0.9098 -0.2382 -0.2875 2.7043 2.9186
0.5 1.0152 0.9533 -0.2388 -0.3123 2.5054 2.8533
0 0 1.0000 1.$0$0.4002 -0.4002 2.5981 2.5981
0.3 1.0518 1.0220 -0.3965 -0.4124 2.4294 2.5658
0.4 1.1232 1.0609 -0.4136 -0.4408 2.2730 2.5176
0.5 1.4142 1.1714 -0.7332 -0.5677 2.0000 2.4231
0.2 0 1.2209 1.2209 -0.7700 -0.7700 2.1686 2.1686
0.3 1.3717 1.2802 -0.9488 -0.8508 1.9534 2.1200
0.4 1.6922 1.4096 -1.4988 -1.0636 1.7232 2.0413
0.5 - - - - - -
Table 1: The corresponding values of the lensing coefficients under specific spin parameters and hair parameters. When $a=0$ and $Q_{m}=0$, the BH degenerates into a Schwarzschild BH. When $a=0$ and $Q_{m}\neq 0$, the BH becomes a Reissner-Nordström (R-N) BH. When $a\neq 0$ and $Q_{m}=0$, the BH is a Kerr BH. The parameter $k=1$ corresponds to the Kerr-Newman BH, and $k=1.5$ corresponds to the rotating short-haired BH.
Figure 6: The variation of the deflection angle $\alpha_{D}$ of a rotating short-haired BH with the impact parameter $u$ under the strong deflection limit. From left to right, the spin parameters are $a=-0.2$ and $a=0.2$.
@@ IV.1 Characteristic Observation Values and Time Delays in Strong Lensing Effects
In gravitational lensing, if the positions of the lens and the light source are known, the specific positions of the images can be calculated using the lens equation. The lens equation is given in Virbhadra and Ellis (2000), and later approximated using the small-angle approximation Bozza (2008b); Bozza et al. (2001)
@@
To approximate the deflection $\Delta\alpha_{n}$, we use the analysis from Bozza (2003). For the $n$-th image, the relationship is given by
@@
Here, $\theta_{n}^{0}$ corresponds to the angle $\theta$ when $\alpha=2n\pi$. By combining these relationships, the approximate position of the $n$-th image can be obtained as Bozza (2002)
@@
Of course, in a lensing system, apart from the corresponding angular positions, the magnification is also an observable physical quantity of great significance. The magnification of the $n$th image can be expressed as Bozza (2002); Virbhadra et al. (1998); Virbhadra and Keeton (2008)
@@
It is evident from the above expression that due to the presence of the $e_{n}$ term, the magnification of the image decreases exponentially with increasing $n$. Additionally, when $\beta$ is very small and approaches zero, we can obtain relatively bright images. In other words, images are easier to observe when $\beta\to 0$. Clearly, the magnification is largest when $n=1$, meaning the image is the brightest at this point. Therefore, if we resolve the first-order image $\theta_{1}$ (the dominant image) and represent the other unresolved images as $\theta_{\infty}$, we can obtain several interesting observables Bozza (2002)
@@
In addition to the observational effects mentioned above, time delay is also an important physical quantity. In the study of gravitational lensing effects, when light passes by a BH, its path bends due to gravity, forming multiple images. These images, because of different light paths, reach the observer at different times, which is known as time delay. This phenomenon can be used to determine the geometric scale and mass of the lensing system in observations and to estimate the Hubble constant in a cosmological context, such as Birrer et al. (2024); Treu et al. (2022); Grillo et al. (2018).Here, we use the method proposed by Bozza and Mancini for calculating time delay in the strong field limit Bozza and Mancini (2004). For a lensing system aligned on a straight line ($\beta=0$) and two relativistic images on the same side of the lens (the $p$-th and $q$-th images), the time delay formula is given by Bozza and Mancini (2004)
@@ IV.2 Evaluating the Observability of Supermassive Black Holes
For the three interesting observables, when regarding the rotating short-haired BH as a candidate for M87* and Sgr A*, we carried out numerical calculations and presented the changes of these observables, as depicted in Figure 7. We selected the case of the rotating short-haired BH corresponding to the parameter $k=1.5$, and calculated the angular position $\theta_{\infty}$ of the relativistic image, the angular distance $s$, and the image magnification $r_{\text{mag}}$. The results show that, whether for Sgr A* or M87*, the variation trends of the observables are identical. Specifically, the angular position $\theta_{\infty}$ of the relativistic image decreases as the spin parameter $a$ and the hair parameter $Q_{m}$ increase. The angular distance $s$ increases with the increase of $a$ and $Q_{m}$. The image magnification $r_{\mathrm{mag}}$ gradually decreases as $a$ increases. In particular, as $Q_{m}$ increases, the variation trend becomes more significant (this is reflected by the density of the contour lines in the figure). By combining the data in Table 2, it can be seen that for M87*, the range of influence of the hair parameter $Q_{m}$ on the angular position of the relativistic image is $22.7113\ \mu as>\theta_{\infty}(M87*)>13.9162\ \mu as$. This range lies within the measurement range of $(42\pm 3)\ \mu as$ when the EHT observes the diameter of M87* Akiyama et al. (2019a). Regarding Sgr A*, the range of influence of the hair parameter is $29.3132\ \mu as>\theta_{\infty}(SgrA*)>17.9614\ \mu as$, and this range coincides with the measured value of $(51.8\pm 2.3)\ \mu as$ for the diameter of the Sgr A* shadow observed by the EHT Akiyama et al. (2022a). Obviously, these observed values are all at the $\mu as$ level, matching the observation accuracy of the existing EHT. Regrettably, though, the resolution of the current equipment is insufficient to detect these differences.
@@
As can be seen from Table 3, under the same spin parameter conditions, for M87*, the deviations in observables between the rotating short-haired BH and the Kerr BH ($Q_{m}=0$) are $\delta\theta_{\infty}\approx 3\ \mu as$, $\delta S\approx 0.1272\ \mu as$, $\delta r_{mag}\approx 0.8980$; and the deviations between the rotating short-haired BH and the Kerr-Newman black hole ($k=1$) are $\delta\theta_{\infty}\approx 2.4249\ \mu as$, $\delta S\approx 0.0535\ \mu as$, $\delta S_{mag}\approx 1.1001$. For Sgr A*, the deviations between the rotating short-haired BH and the Kerr BH are $\delta\theta_{\infty}\approx 3.3786\ \mu as$, $\delta S\approx 0.1642\ \mu as$, $\delta r_{mag}\approx 0.8980$; and the deviations between the rotating short-haired BH and the Kerr-Newman BH are $\delta\theta_{\infty}\approx 3.1299\ \mu as$, $\delta S\approx 0.0690\ \mu as$, $\delta S_{mag}\approx 1.1001$. Although these deviations are all at the $\mu as$ level (especially the angular position), the angular resolution of the current EHT is approximately $20\ \mu as$, so it is impossible to distinguish these small differences Akiyama et al. (2019c). Therefore, to distinguish between the rotating short-haired BH and the Kerr BH or the Kerr-Newman BH may depend on the next-generation EHT. Once these differences can be distinguished in the future, it will provide us with further information about the short-hair parameter $Q_{m}$, and also offer an opportunity to test the no-hair theorem.
@@
Table 3: When considering M87* and Sgr A* as the rotating short-haired BH, the deviations in the observed values of the rotating short-haired BH from those in the cases of Kerr BH and Kerr-Newman BH. Here, $\delta(X)=X_{\mathrm{hair}}-X_{\mathrm{Kerr}}$ or $\delta(X)=X_{\mathrm{hair}}-X_{\mathrm{KN}}$.
@@
$k=1$ $k=1.5$ $k=1$ $k=1.5$
$a$ $Q_{m}$ $\Delta T_{21}(h)$ $\Delta T_{21}(h)$ $\Delta T_{21}(min)$ $\Delta T_{21}(min)$
-0.2 0 332.1074 332.1074 12.2011 12.2011
0.3 315.8604 329.3723 11.6042 12.1006
0.4 301.5053 325.3982 11.0768 11.9546
0.5 279.3262 318.1182 10.2620 11.6872
0 0 289.6618 289.6618 10.6417 10.6417
0.3 270.8501 286.0676 9.9506 10.5097
0.4 253.4180 280.6904 9.3102 10.3121
0.5 222.9818 270.1475 8.1920 9.9248
0.2 0 241.7744 241.7744 8.8824 8.8824
0.3 217.7821 236.3585 8.0010 8.6834
0.4 192.1213 227.5810 7.0582 8.3610
0.5 - 203.4960 - 7.4761
@@
Figure 8: When considering M87* and Sgr A* as the rotating short-haired BH, the time delay deviation $\delta\Delta T_{21}$ of the same image between the rotating short-haired BH and the Kerr BH. Here, $\delta(X)=X_{\mathrm{hair}}-X_{\mathrm{Kerr}}$.
@@
Figure 9: When M87* and Sgr A* are regarded as the rotating short-haired BH, the time delay deviation of the same image, $\delta\Delta T_{21}$, between the rotating short-haired BH and the Kerr BH. Here, $\delta(X)=X_{\mathrm{hairy}}-X_{K-N}$.
@@
For the two relativistic images located on the same side of the BH, the time delay between the second relativistic image ($p=2$) and the first relativistic image ($q=1$) is presented in Table 4. The deviations between the rotating short-haired BH, the Kerr BH, and the Kerr-Newman BH are shown in Table 4 as well as Figures 8 and 9. Here, the time delay we consider is mainly determined by the optical path difference between relativistic images formed after light rays orbit the BH several times (see Figure 6. Obviously, near the divergence of the deflection angle, the light rays orbit more than $2\pi$). When considering M87* and SgrA* as a rotating short-haired BH, the time delay $\mathrm{\Delta}T_{2,1}$ between the second and the first relativistic images on the same side is such that the time delay of the former can be up to several hundred hours, while that of the latter can reach tens of minutes. Evidently, the time delay of the former is sufficient for astronomical observations, which provides a necessary condition for exploring the properties of the short-haired BH (see Table 4). Under the same circumstances, the time delay deviation amounts between the rotating short-haired BH on one hand, and the Kerr BH and the Kerr-Newman BH on the other hand: In the case of Sgr A* as the background, the time delay deviation between the rotating short-haired BH and the Kerr BH can reach up to more than two minutes (see Figure 8), while the time delay deviation between the rotating short-haired BH and the Kerr-Newman BH can reach up to more than four minutes (see Figure 9). In the case of M87* as the background, the time delay deviation between the rotating short-haired BH and the Kerr BH can reach up to more than fifty hours (see Figure 8), and the time delay deviation between the rotating short-haired BH and the Kerr-Newman BH can reach up to more than one hundred hours (see Figure 9). Overall, for Sgr A*, the time delay deviation is in the range of several minutes, which is difficult to be detected in astronomy. For M87*, the time delay can reach hundreds of hours. The time delay deviation between it (when considering M87* as a rotating short-haired BH) and the Kerr BH can reach over fifty hours, and the deviation between it and the Kerr-Newman BH can reach over one hundred hours. These time durations are sufficient for astronomical observations. Just from the perspective of time delay, if M87* is regarded as a candidate for a rotating short-haired BH, it is feasible to distinguish between traditional BHs and rotating short-haired BHs. However, the prerequisite for this is that the equipment used is capable of resolving the two relativistic images. Obviously, current equipment fails to meet this requirement. But the next generation EHT is expected to obtain such resolution. If these images can be distinguished in the near future, a lot of important physical information will be obtained, such as testing the no-hair theorem, exploring the properties of the hair parameter $Q_{m}$, and restricting the variation of parameters.

### v3

@@ III Strong Gravitational Lensing in the Spacetime of the Rotating Short-haired Black Hole
$a$ $Q_{m}$ $\bar{a}$ $\bar{b}$ $u_{m}/R_{s}$ $u_{m}/R_{s}$ $u_{m}/R_{s}$
$k=1$ $k=1.5$ $k=2$ $k=1$ $k=1.5$ $k=2$ $k=1$ $k=1.5$ $k=2$
-0.2 0 0.8796 0.8796 0.8796 -0.2810 -0.2810 -0.2810 2.9788 2.9788 2.9788
0.3 0.9048 0.8911 0.8833 -0.2570 -0.2824 -0.2838 2.8331 2.9543 2.9746
0.4 0.9354 0.9098 0.8917 -0.2382 -0.2875 -0.2907 2.7043 2.9186 2.9653
0.5 1.0152 0.9533 0.9123 -0.2388 -0.3123 -0.3106 2.5054 2.8533 2.9446
0 0 1.0000 1.0000 1.$0$0.4002 -0.4002 -0.4002 2.5981 2.5981 2.5981
0.3 1.0518 1.0220 1.0074 -0.3965 -0.4124 -0.4048 2.4294 2.5658 2.5918
0.4 1.1232 1.0609 1.0252 -0.4136 -0.4408 -0.4280 2.2730 2.5176 2.5777
0.5 1.4142 1.1714 1.0733 -0.7332 -0.5677 -0.4920 2.0000 2.4231 2.5452
0.2 0 1.2209 1.2209 1.2209 -0.7700 -0.7700 -0.7700 2.1686 2.1686 2.1686
0.3 1.3717 1.2802 1.2423 -0.9488 -0.8508 -0.8051 1.9534 2.1200 2.1573
0.4 1.6922 1.4096 1.2989 -1.4988 -1.0636 -0.9060 1.7232 2.0413 2.1310
0.5 - - - - - - - - -
Table 1: The values corresponding to the lensing coefficients under specific spin parameters and hair parameters. When $a=0$ and $Q_{m}=0$, the BH degenerates to a Schwarzschild BH; when $a=0$ and $Q_{m}\neq 0$, the BH becomes a Reissner-Nordström BH; when $a\neq 0$ and $Q_{m}=0$, the BH is a Kerr BH. The parameter $k=1$ corresponds to the Kerr-Newman BH, while $k=1.5$ and $k=2$ correspond to rotating short-haired BH solutions under different parameter configurations.
Figure 6: Deflection angle characteristics of rotating short-haired BHs in the strong deflection limit. (a) The variation of deflection angle with impact parameter $u$ under different hair parameters, with spin parameter $a=-0.2$ and hair parameter $k=1.5$; (b) The variation of deflection angle with impact parameter $u$ under different hair parameters, with spin parameter $a=0.2$ and hair parameter $k=1.5$; (c) The relationship between deflection angle and hair parameter $k$ for different hair parameter values, with spin parameter $a=0.2$ and impact parameter $u=2.5$. The black dashed line represents the Kerr BH case.
@@ IV.1 Characteristic Observation Values and Time Delays in Strong Lensing Effects
In gravitational lensing, if the positions of the lens and the light source are known, the specific positions of the images can be calculated using the lens equation. The lens equation is given in [33], and later approximated using the small-angle approximation [59, 31]
@@
To approximate the deflection $\Delta\alpha_{n}$, we use the analysis from [32]. For the $n$-th image, the relationship is given by
@@
Here, $\theta_{n}^{0}$ corresponds to the angle $\theta$ when $\alpha=2n\pi$. By combining these relationships, the approximate position of the $n$-th image can be obtained as [34]
@@
Of course, in a lensing system, apart from the corresponding angular positions, the magnification is also an observable physical quantity of great significance. The magnification of the $n$th image can be expressed as [34, 60, 61]
@@
It is evident from the above expression that due to the presence of the $e_{n}$ term, the magnification of the image decreases exponentially with increasing $n$. Additionally, when $\beta$ is very small and approaches zero, we can obtain relatively bright images. In other words, images are easier to observe when $\beta\to 0$. Clearly, the magnification is largest when $n=1$, meaning the image is the brightest at this point. Therefore, if we resolve the first-order image $\theta_{1}$ (the dominant image) and represent the other unresolved images as $\theta_{\infty}$, we can obtain several interesting observables [34]
@@
In addition to the observational effects mentioned above, time delay is also an important physical quantity. In the study of gravitational lensing effects, when light passes by a BH, its path bends due to gravity, forming multiple images. These images, because of different light paths, reach the observer at different times, which is known as time delay. This phenomenon can be used to determine the geometric scale and mass of the lensing system in observations and to estimate the Hubble constant in a cosmological context, such as [62, 63, 64].Here, we use the method proposed by Bozza and Mancini for calculating time delay in the strong field limit [65]. For a lensing system aligned on a straight line ($\beta=0$) and two relativistic images on the same side of the lens (the $p$-th and $q$-th images), the time delay formula is given by [65]
@@ IV.2 Evaluating the Observability of Supermassive Black Holes
For the three interesting observables, when regarding the rotating short-haired BH as a candidate for M87* and Sgr A*, we carried out numerical calculations and presented the changes of these observables. As illustrated in Figure 7, we selected the case of a rotating short-haired BH corresponding to parameter $k=1.5$, and calculated the angular position of relativistic images $\theta_{\infty}$, angular separation $S$, and image magnification ratio $r_{\text{mag}}$ under different hair parameters $Q_{m}$. The results show that, whether for Sgr A* or M87*, the variation trends of the observables are identical. Specifically, the angular position $\theta_{\infty}$ of the relativistic image decreases as the spin parameter $a$ and the hair parameter $Q_{m}$ increase. The angular separation $S$ increases with the increase of $a$ and $Q_{m}$. The image magnification ratio $r_{\mathrm{mag}}$ gradually decreases as $a$ increases. In particular, as $Q_{m}$ increases, the variation trend becomes more significant (this is reflected by the density of the contour lines in the figure). By combining the data in Table 2, it can be seen that for M87*, the range of influence of the hair parameter $Q_{m}$ on the angular position of the relativistic image is $22.7113\ \mu as>\theta_{\infty}(M87*)>13.9162\ \mu as$. This range lies within the measurement range of $(42\pm 3)\ \mu as$ when the EHT observes the diameter of M87* [48]. Regarding Sgr A*, the range of influence of the hair parameter is $29.3132\ \mu as>\theta_{\infty}(SgrA*)>17.9614\ \mu as$, and this range coincides with the measured value of $(51.8\pm 2.3)\ \mu as$ for the diameter of the Sgr A* shadow observed by the EHT [49]. Obviously, these observed values are all at the $\mu as$ level, matching the observation accuracy of the existing EHT. Regrettably, though, the resolution of the current equipment is insufficient to detect these differences.
@@
As can be seen from Table 4, under the same spin parameter conditions, for M87*, the deviations in observables between the rotating short-haired BH and the Kerr BH ($Q_{m}=0$) are $\delta\theta_{\infty}\approx 3\ \mu as$, $\delta S\approx 0.1272\ \mu as$, $\delta r_{mag}\approx 0.8980$; and the deviations between the rotating short-haired BH and the Kerr-Newman black hole ($k=1$) are $\delta\theta_{\infty}\approx 2.4249\ \mu as$, $\delta S\approx 0.0535\ \mu as$, $\delta S_{mag}\approx 1.1001$. For Sgr A*, the deviations between the rotating short-haired BH and the Kerr BH are $\delta\theta_{\infty}\approx 3.3786\ \mu as$, $\delta S\approx 0.1642\ \mu as$, $\delta r_{mag}\approx 0.8980$; and the deviations between the rotating short-haired BH and the Kerr-Newman BH are $\delta\theta_{\infty}\approx 3.1299\ \mu as$, $\delta S\approx 0.0690\ \mu as$, $\delta S_{mag}\approx 1.1001$. Although these deviations are all at the $\mu as$ level (especially the angular position), the angular resolution of the current EHT is approximately $20\ \mu as$, so it is impossible to distinguish these small differences [68]. Therefore, to distinguish between the rotating short-haired BH and the Kerr BH or the Kerr-Newman BH may depend on the next-generation EHT. Once these differences can be distinguished in the future, it will provide us with further information about the short-hair parameter $Q_{m}$, and also offer an opportunity to test the no-hair theorem.
In addition, from Figure 8 and Table 3, it can be observed that as the parameter $k$ increases, all corresponding observables gradually approach the case of the standard Kerr BH (the red dashed line in Figure a1 and the values shown in the header of Table 3 represent the Kerr BH case with $Q_{m}=0$). This result indicates that larger $k$ values actually suppress the influence of the hair parameter, causing the differences in observational characteristics between rotating short-haired BH and Kerr BH to gradually diminish to an indistinguishable degree. Based on this consideration, we chose to fix the parameter $k=1.5$ in the above analysis, a choice that both maintains the rationality of the physical model and adequately demonstrates the characteristic influence of the hair parameter on observables. This allows us to effectively quantify and evaluate the contributions of hair parameters and spin to various observables.
@@
Figure 8: When considering M87* and Sgr A* as a rotating short-haired BH, the variation of angular position $\theta_{\infty}$ with parameter $k$ under different hair parameters. The top row is for M87*, and the bottom row is for Sgr A*. The left column corresponds to the case where $a=0$, and the right column corresponds to the case where $a=0.2$. The red dotted line in the figure represents the case of a Kerr BH.
@@
$Q_{m}=0.1$ $Q_{m}=0.3$ $Q_{m}=0.1$ $Q_{m}=0.3$ $Q_{m}=0.1$ $Q_{m}=0.3$
$a$ $k$ $\theta_{\infty}(\mu as)$ $S(\mu as)$ $\theta_{\infty}(\mu as)$ $S(\mu as)$ $\theta_{\infty}(\mu as)$ $S(\mu as)$ $\theta_{\infty}(\mu as)$ $S(\mu as)$ $r_{mag}$ $r_{mag}$
0 1 19.6756 0.0254 18.5222 0.0323 25.3950 0.0328 23.9064 0.0417 6.7909 6.4857
1.5 19.7998 0.0249 19.5629 0.0279 25.5554 0.0322 25.2495 0.0361 6.8168 6.6750
2 19.8081 0.0248 19.7606 0.0258 25.5660 0.0320 25.5047 0.0333 6.8213 6.7716
2.5 19.8086 0.0248 19.7991 0.0251 25.5667 0.0320 25.5545 0.0324 6.8218 6.8065
3 19.8087 0.0248 19.8068 0.0249 25.5668 0.0320 25.5643 0.0321 6.8219 6.8175
0.2 1 16.3708 0.0531 14.8931 0.0764 21.1295 0.0685 19.2224 0.0986 5.5342 4.9733
1.5 16.5209 0.0515 16.1635 0.0614 21.3233 0.0665 20.8620 0.0793 5.5788 5.3286
2 16.5328 0.0513 16.4479 0.0547 21.3387 0.0662 21.2291 0.0706 5.5862 5.4914
2.5 16.5338 0.0512 16.5135 0.0524 21.3399 0.0661 21.3137 0.0676 5.5872 5.5541
3 16.5339 0.0512 16.5290 0.0516 21.3400 0.0661 21.3337 0.0666 5.5874 5.5764
Table 3: The observable values corresponding to different parameters $k$ under fixed hair parameters $Q_{m}=0.1$ or $Q_{m}=0.3$ when considering M87* and Sgr A* as a rotating short-haired BH. For comparison reference, the standard Kerr BH $(Q_{m}=0)$ at spin parameter $a=0$ has corresponding observable values of $\theta_{\infty}=19.8087\,\mu as$, $S=0.0248\,\mu as$, $r=6.8219\,\mu as$ (for M87*), $\theta_{\infty}=25.5668\,\mu as$, $S=0.0320\,\mu as$, $r=6.8219\,\mu as$ (for Sgr A*); at spin parameter $a=0.2$, its observable parameters are $\theta_{\infty}=16.5339\,\mu as$, $S=0.0512\,\mu as$, $r=6$ (for M87*), $\theta_{\infty}=21.3400\,\mu as$, $S=0.0661\,\mu as$, $r=5.5874$ (for Sgr A*).
$M87^{*}$ $SgrA^{*}$ $M87^{*}orSgrA^{*}$
@@
Table 4: When considering M87* and Sgr A* as the rotating short-haired BH, the deviations in the observed values of the rotating short-haired BH from those in the cases of Kerr BH and Kerr-Newman BH. Here, $\delta(X)=X_{\mathrm{hair}}^{k=1.5}-X_{\mathrm{Kerr}}$ or $\delta(X)=X_{\mathrm{hair}}^{k=1.5}-X_{\mathrm{KN}}$.
@@
$k=1$ $k=1.5$ $k=2$ $k=1$ $k=1.5$ $k=2$
$a$ $Q_{m}$ $\Delta T_{21}(h)$ $\Delta T_{21}(h)$ $\Delta T_{21}(h)$ $\Delta T_{21}(min)$ $\Delta T_{21}(min)$ $\Delta T_{21}(min)$
-0.2 0 332.1074 332.1074 332.1074 12.2011 12.2011 12.2011
0.3 315.8604 329.3723 331.637 11.6042 12.1006 12.1838
0.4 301.5053 325.3982 330.5983 11.0768 11.9546 12.1457
0.5 279.3262 318.1182 328.2968 10.2620 11.6872 12.0611
0 0 289.6618 289.6618 289.6618 10.6417 10.6417 10.6417
0.3 270.8501 286.0676 288.959 9.9506 10.5097 10.6159
0.4 253.4180 280.6904 287.3847 9.3102 10.3121 10.5581
0.5 222.9818 270.1475 283.7672 8.1920 9.9248 10.4252
0.2 0 241.7744 241.7744 241.7744 8.8824 8.8824 8.8824
0.3 217.7821 236.3585 240.5179 8.0010 8.6834 8.8363
0.4 192.1213 227.5810 237.5862 7.0582 8.3610 8.7286
0.5 - 203.4960 229.9303 - 7.4761 8.4473
@@
Figure 9: When considering M87* and Sgr A* as the rotating short-haired BH, the time delay deviation $\delta\Delta T_{21}$ of the same image between the rotating short-haired BH and the Kerr BH. Here, $\delta(X)=X_{\mathrm{hair}}^{k=1.5}-X_{\mathrm{Kerr}}$.
@@
Figure 10: When M87* and Sgr A* are regarded as the rotating short-haired BH, the time delay deviation of the same image, $\delta\Delta T_{21}$, between the rotating short-haired BH and the Kerr BH. Here, $\delta(X)=X_{\mathrm{hairy}}^{k=1.5}-X_{\text{K-N}}$.
@@
For the two relativistic images located on the same side of the BH, the time delay between the second relativistic image ($p=2$) and the first relativistic image ($q=1$) is presented in Table 5. The deviations between the rotating short-haired BH ($k=1.5$), the Kerr BH, and the Kerr-Newman BH are shown in Table 5 as well as Figures 9 and 10. Here, the time delay we consider is mainly determined by the optical path difference between relativistic images formed after light rays orbit the BH several times (see Figure 6. Obviously, near the divergence of the deflection angle, the light rays orbit more than $2\pi$). When considering M87* and SgrA* as a rotating short-haired BH, the time delay $\mathrm{\Delta}T_{2,1}$ between the second and the first relativistic images on the same side is such that the time delay of the former can be up to several hundred hours, while that of the latter can reach tens of minutes. Evidently, the time delay of the former is sufficient for astronomical observations, which provides a necessary condition for exploring the properties of the short-haired BH (see Table 5). Under the same circumstances, the time delay deviation amounts between the rotating short-haired BH on one hand, and the Kerr BH and the Kerr-Newman BH on the other hand: In the case of Sgr A* as the background, the time delay deviation between the rotating short-haired BH and the Kerr BH can reach up to more than two minutes (see Figure 9), while the time delay deviation between the rotating short-haired BH and the Kerr-Newman BH can reach up to more than four minutes (see Figure 10). In the case of M87* as the background, the time delay deviation between the rotating short-haired BH and the Kerr BH can reach up to more than fifty hours (see Figure 9), and the time delay deviation between the rotating short-haired BH and the Kerr-Newman BH can reach up to more than one hundred hours (see Figure 10). Overall, for Sgr A*, the time delay deviation is in the range of several minutes, which is difficult to be detected in astronomy. For M87*, the time delay can reach hundreds of hours. The time delay deviation between it (when considering M87* as a rotating short-haired BH) and the Kerr BH can reach over fifty hours, and the deviation between it and the Kerr-Newman BH can reach over one hundred hours. These time durations are sufficient for astronomical observations. Just from the perspective of time delay, if M87* is regarded as a candidate for a rotating short-haired BH, it is feasible to distinguish between traditional BHs and rotating short-haired BHs. However, the prerequisite for this is that the equipment used is capable of resolving the two relativistic images. Obviously, current equipment fails to meet this requirement. But the next generation EHT is expected to obtain such resolution. If these images can be distinguished in the near future, a lot of important physical information will be obtained, such as testing the no-hair theorem, exploring the properties of the hair parameter $Q_{m}$, and restricting the variation of parameters.